OpenStudy (anonymous):

How many pairs of natural numbers are there so that difference of their squares is 60

OpenStudy (anonymous):

i got only 2 pairs of solutions work, a = 8, b = 2, and a = 16, b = 14

OpenStudy (anonymous):

going to post a pick of my solution in a sec.

OpenStudy (anonymous):

good morning joemath up all night?

OpenStudy (anonymous):

Alright, this solution revolves around that fact that: \[a^{2} - b^{2} = (a+b)(a-b)\] Because we are dealing with natural numbers, a+b and a-b are also natural numbers. So we have 2 natural numbers that multiply to 60, we means they must be factors of 60.

OpenStudy (anonymous):

morning satelitte :) i dont know if you can call it up all night, i work up around 3 am ish lol

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yikes. this problem is asking how many pythgorean triples there are with one number = 60

OpenStudy (anonymous):

just thought i would mention it

OpenStudy (anonymous):

So basically what i did was check the cases. say if a + b = 10, then a-b would have to be 6, and that leads to a valid solution. infact, the only valid solutions are a+b = 10, a-b = 6, and a+b = 30, a- b = 2, so there are 2 pairs of natural numbers that solve that equation.

OpenStudy (anonymous):

how did you scan so kwikly

OpenStudy (anonymous):

funny story: i bought a new scanner so I could beat satellite to solutions :P (true story!)

OpenStudy (anonymous):

cool